OpenStudy (anonymous):

Suppose the square matrix C={1 1 -1} 3 0 -1 -1 -1 2 1) Evaluate det(C) by expanding along the first row, Include steps of the calculation. NO MARKS awarded if any other method is used.

OpenStudy (anonymous):

Do you know how to calculate determinant ??

OpenStudy (anonymous):

Yes I do, but I don't understand what the expand along the first row means.

OpenStudy (anonymous):

Using first row... Tell me how you calculate determinant?? Calculate the determinant for this by what ever the way you know ..

OpenStudy (anonymous):

Using first row means you will use : 1 1 and -1 for calculating the determinant.. I will show you in general what do you get using first row for determinant ..

OpenStudy (anonymous):



OpenStudy (anonymous):

I'm sorry Waterineyes, but I don't understand why you minused the b part of the equation, was that an error or it's the way it is.

OpenStudy (anonymous):

No it is not the error..

OpenStudy (anonymous):

Generally there are sign given to all the positions of matrix..

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh thanks a lot, I think I understand now, let me try doing the exercise... Will be back in 2 minutes...

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

One more thing, do we consider all the signs when doing the exercise? because I've got 2 different answers, in 1 I didn't consider the signs and the second 1 I considered the signs. I then got 9 and 8.

OpenStudy (anonymous):

Can you show how you got them so that I can find out the mistake or I can see them with my eyes..??

OpenStudy (anonymous):

No, when expanding along first row, you will use the signs of first row only.. For example : a and c will be positive and b will be negative.. Do not use the signs for other values like d e f g h i.. Leave these values as such... Use the signs for a b and c only..

OpenStudy (anonymous):

See here : Det(C) = \(1(0 \cdot2 - (-1) \cdot (-1) - 1 (3 \cdot 2 - (-1) \cdot (-1) + -1(3 \cdot (-1) - (-1) \cdot 0))\)

OpenStudy (anonymous):

Simplifying little bit you will get like this : Det(C) = \(1(-1) - 1(5) - 1(-3)\) Solve this to get the determinant..